OpenStudy (anonymous):

Can someone please show me the rest of the steps to solve this problem? Write the equation in standard form to find the center and the radius of the circle. x^2-5=-y^2-6y +y^2+6y +y^2 +6y -------------------- x^2-5+y^2+6y=0 +5 +5 -------------------- x^2+y^2+6y=5 (x^2____________) + (y^2+6y+9)=5+9 There's no coefficient for the variable x so I'm not sure what I'm supposed to do with the x^2?

Directrix (directrix):

The coefficient of the linear x variable is 0 as in 0*x. Hence, no linear x term appears. Your work looks good. (x^2____________) + (y^2+6y+9)=5+9 (x^2 + 0x + 0 )+ (y^2+6y+9)=5+9 (x + 0) ^2 + ( y + ? )^2 = 14 Finish by writing equation in standard form. See attached chart if a refresher for standard form is needed.